# AI-based Programming Assistants for Privacy-related Code Generation: The Developers’ Experience

### v1

@@
Abstract
With the popularising of generative AI, the existence of AI-based programming assistants for developers is no surprise. Developers increasingly use them for their work, including generating code to fulfil the data protection requirements (privacy) of the apps they build. We wanted to know if the reality is the same as expectations of AI-based programming assistants when trying to fulfil software privacy requirements, and the challenges developers face when using AI-based programming assistants and how these can be improved. To this end, we conducted a survey with 51 developers worldwide. We found that AI-based programming assistants need to be improved in order for developers to better trust them with generating code that ensures privacy. In this paper, we provide some practical recommendations for developers to consider following when using AI-based programming assistants for privacy-related code development, and some key further research directions.
Index Terms:
@@ Abstract
I Introduction
Neglecting implementing privacy requirements or taking privacy lightly could lead to disastrous consequences, including catastrophic data breaches [1]. Attackers may obtain sensitive data of end users – usernames, passwords, banking details, location, addresses etc., – with malicious intentions unless the software is highly secure from diverse data breaches [2]. In September 2022, in Australia, a major data breach happened on Optus (a telecommunication service) where passport details, driving license numbers, and Medicare (health care insurance scheme in Australia) information of over 10 million current and former customers were hacked [3]. In April 2021, one of the most controversial and largest data breaches happened on Facebook where names, phone numbers, usernames, and passwords of 530 million users were exposed [4]. The biggest data breach in US history happened on Yahoo from 2013 to 2016 where personally identifiable information of over 3 billion user accounts was exposed [5]. These are only a few of many examples of dreadful data breaches. Developers must thus focus on better-protecting user data, including code logic, interfaces, and testing [6].
@@ I Introduction
Given the capacity of generating human-like text using the large corpus of data the large language models (LLMs) were trained on, the popularity of using LLMs in applications of different domains is rising exponentially. Codex – OpenAI’s LLM model used to generate code in GitHub’s Copilot – claims that it is ”proficient in more than a dozen programming languages, Codex can now interpret simple commands in natural language and execute them on the user’s behalf—making it possible to build a natural language interface to existing applications” [^1] https://openai.com/index/openai-codex/. Similar to Copilot, other LLM-driven code generators (AI-based programming assistants) such as Amazon Q Developer (Amazon Code Whisperer), Tabnine, Kite etc., are widely available for the developer community to use. These AI-based programming assistants claim that they enhance productivity, which is key in software development [^2] https://aws.amazon.com/q/developer/, https://github.com/features/copilot and some research has found the same results [7, 8]. However, AI-based programming assistants struggle with complex tasks [9], the accuracy of the code of these AI-based programming assistants depends on the programming language and the task [10, 11, 12], with longer details resulting in poor results [13], and developers may spend more time reviewing the code that AI-based programming assistants produced [14]. The developers also face challenges with understanding, editing, and debugging the code generated by the tools [15, 16], resulting in inefficiency [17].
GDPR defines data protection (“privacy” as used in this paper), as “keeping data safe from unauthorised access” [18]. Privacy engineering encourages embedding privacy into systems. For example, designing clear privacy controls (on the user-facing side), determining the best methods for anonymisation, and inspecting code before deployment to assess privacy risks. Proactive measures of privacy engineering include data minimisation and retention [19]. One such proactive technique is privacy by design. The implementation of privacy by design varies by systems used, the project undertaken, and the extent to which the personal information is dealt. The principles of privacy by design are proactive, not reactive, privacy as the default setting, privacy embedded into design, full functionality, end-to-end security, visibility and transparency, and respect for user privacy. For a business to leverage the data it has while ensuring the privacy of its users’ personal data, privacy-enhancing technologies (PETs) can be used. PETs are important due to the necessity of following data protection laws such as GDPR and CCPA, requirements of testing data by third parties, and protecting the business from tarnishing its reputation from privacy breaches. PETs such as cryptographic algorithms (e.g., homomorphic encryption, secure multi-party computation), data masking techniques (e.g., obfuscation, pseudonymization), and AI and machine learning algorithms (e.g., synthetic data generation, federated learning) exist for developers to use.
@@
However, developers struggle with embedding privacy into software systems [20, 21, 22]. They find it difficult to relate privacy requirements to privacy techniques, they see privacy in contradiction with system requirements, developers personal and peers’ opinions affect how they design, and they lack privacy knowledge [23]. When developers are asked to consider privacy in application designs, developers intuitively focus on technical aspects, but not user privacy expectations. Developers are also observed to have a reduced level of privacy expectations compared to users [24]. This leads to several critical issues. For example, permission-related issues frequently happen in Android apps [25]. Having tools such as AI-based programming assistants comes in handy for developers in all the above-mentioned cases. Therefore, the developers may want to use AI-based programming assistants to generate code to meet their software’s privacy requirements.
@@
Nevertheless, the widespread uptake of AI-based programming assistants to implement systems – including data privacy-related code – raises the potential of their use for more secure code construction but, on the other hand, also introduces unexpected and unintended code vulnerabilities [26]. To better understand the practical experience of using AI-based programming assistants for privacy-related code generation, we carried out a survey [^3] Approved by Monash Human Research Ethics Committee. Project ID: 37904. of 51 software developers worldwide. We asked the developers about their experience with AI-based programming assistants for privacy-related code generation. Our research yielded knowledge and insights on AI-based programming assistants including the difference between the expectation and reality of AI-based programming vis-à-vis meeting privacy-related requirements in AI-generated code and the critical need for improvements in the AI-based programming assistants for developers to trust them with generating code for ensuring privacy.
II Research Questions
Consider Xiaoyu, a developer who wants to make their programming life better and more importantly wants to ensure they write accurate code to meet privacy requirements which are fundamental and crucial for the software she is developing. They want to use an off-the-shelf AI-based programming assistant to assist them in generating privacy-preserving code. They need and expect the AI to generate quality privacy-preserving code - but does it? What setbacks do they encounter? What can improve the AI-based programming assistants to generate privacy-preserving code? We wanted to answer the following research questions formulated from such a scenario:
@@ II Research Questions
    Is the reality the same as the expectation of AI-based programming assistants when it comes to ensuring privacy requirements are met in the code? (Section IV-B)
    Privacy is a crucial non-functional requirement in software development. We deep-dived into this aspect on the subject of the code generated by AI-based programming assistants. As the first step, we evaluated if the reality is similar or different to developers’ expectations.
@@
    What are the challenges with generating AI-based programming assistants to ensure meeting privacy requirements? (Section IV-C)
    Then we wanted to explore the challenges concerning privacy-ensuring code as identifying the challenges is essential for improving the AI-based programming assistants.
@@
    The developers are the users of AI-based programming assistants. It is indispensable to get their feedback on improvements to the AI-based programming assistants based on their experiences with them. We used this question to investigate potential improvements in AI-based programming assistants.
III Our Approach – The Developer Survey
III-A Survey Questionnaire Development
@@
To answer the research questions, we conducted a developer survey. We designed our survey with both open-ended and closed-ended questions with a mapping against our research questions (Table I). For example, to answer RQ1, we used both open-ended and closed-ended questions. For the complete survey questionnaire (this includes questions beyond the scope of this paper), please check our online replication package [^4] https://github.com/kashumi-m/AICodeGensPrivacy. The questions used to collect data on company size, developers’ roles, and employment status reported in this paper were adapted from the Stack Overflow developer survey [^5] https://survey.stackoverflow.co/2022. After we developed the questionnaire, we piloted it with the help of two postdoctoral researchers who worked in the software industry in the past at the HumaniSE Lab at Monash University. They gave feedback on completion time, and the flow of the questions. Upon their feedback, we moved the demographic questions to the end of the questionnaire, which were at the beginning earlier.
TABLE I: We used a mix of open-ended and closed-ended questions to answer the RQs.
RQ Question type
Context Closed-ended questions with an ”other” option to enter text
RQ1 Closed-ended questions and an open-ended question
RQ2 An open-ended question
RQ3 An open-ended question
III-B Data Collection
@@
After finalising the questionnaire, we hosted the survey on Qualtrics and recruited developers who work in the IT industry via the recruitment platform Prolific. We spent 6 Sterling Pounds on each participant. The participant spent around 15-20 minutes to fill out the questionnaire. We recruited 51 participants and we were able to generate rich insights from the data we collected. The details of the participants are given in the next section.
III-C Data Analysis
To analyse the data we collected through open-ended (qualitative) and closed-ended (quantitative) questions, we used open coding and constant comparison [27] – interpreting the data using small chunks of words (codes) and categorising them meaningfully by comparing each other, and descriptive statistics respectively.
IV Findings
IV-A Survey Participants
IV-A1 Demographics of the Participants
The majority of the developers who participated in our study were men (76.47%; n=39) and the participants resided around the world (Fig. 1). Most participants were from companies with 101–500 employees (17.65%; n=9) (Fig. 2). The majority were from the IT industry (78.43%; n=40).
Figure 1: The majority of the developers who participated in the study were Men and resided widely around the world.
Figure 2: Most participants belonged to companies with 101–500 employees. The majority of the participants were from the IT industry.
Figure 3: The majority of our respondents worked full-time, hybrid. They had a median of 9 years of experience in the software industry, 7 years of professional coding experience, and 6 years of agile experience.
@@ IV-A1 Demographics of the Participants
Most of our participants were full-time employees (88.84%; n=45) and the majority worked hybrid – some remote and some in-person (52.94%; n=27) (Fig. 3). The participants had a median of 9 years of experience in the software industry, 7 years of professional coding experience, and 6 years of agile experience.
IV-A2 Who Used AI-based Programming Assistants the Most and Who Used Them the Least
@@
While our participants played multiple roles in their jobs (Fig. 4), the majority were full-stack developers (37.25%; n=19). The majority of the developers used AI-based programming assistants 1–2 times a day (54.90%; n=28). From that, most were full-stack developers (21.57%; n=11). Out of the participants who never used AI-based programming assistants (19.61%; n=10) were the back-end developers (7.84%; n=4). ChatGPT was the most commonly used (61.54%; n=32) AI-based programming assistant by our developers (Fig. 5).
Figure 4: The majority of our participants were full-stack developers and most of them used AI-based programming assistants 1–2 times a day. The most among the developers who never used AI-based programming assistants were back-end developers.
Figure 5: Some developers used more than one AI-based programming assistant. However, ChatGPT is the most commonly used tool.
@@
IV-B Expectation vs Reality of AI-based Programming Assistants for Privacy-related Code Generation (answer to RQ1)
@@
At the first instance of using AI-based programming assistant (tool), the majority of the developers (64.71%; n=33) expected that the tool will generate code ensuring that key data privacy requirements are met (Fig. 6). However, as they moved forward with using the tool, the reality became somewhat different from what they expected. The majority (39.22%; n=20) found that AI-based programming assistants sometimes generate code ensuring privacy requirements without requesting it to meet the requirement (Fig. 7). However, they had to always (52.94%; n=27) put extra effort into double-checking the generated code to see if privacy requirements were met (Fig. 8). Therefore, unsurprisingly, a large number of the developers (43.14%; n=22) never used AI-based programming assistants to generate code for privacy requirements (Fig. 10), and many of them (41.18%, n=21) wrote code manually most of the time to meet privacy requirements (Fig. 9), even though they used AI-based programming assistants during software development.
@@
Figure 6: Developers expected that the AI-based programming assistant will ensure meeting the privacy requirements.
@@
Figure 7: Developers found that AI-based programming assistants sometimes generate code meeting privacy requirements..
@@
Figure 8: Developers always put extra effort to double-check the code generated by the AI tool to ensure the privacy requirements were met.
@@
Figure 9: Most of the time, the majority of the developers wrote code manually to ensure privacy requirements were met.
@@
Figure 10: Majority of the developers never used AI-based programming assistants to generate code to meet privacy requirements.
@@
IV-C Key Challenges with Generating Code Ensuring Privacy using AI-based Programming Assistants (answer to RQ2)
@@
IV-C1 Unfulfilled task - not generating privacy-preserving code (19.61%; n=10)
Our developers mentioned that current AI-based programming assistants lack deep comprehension of and code generation meeting diverse privacy regulations. They said that AI-based programming assistants may not be able to keep up with changing privacy policies, they may not fully understand privacy rules or specific needs. Hence, the AI-based programming assistants may struggle to implement customised privacy controls.
“When I use AI-code generators to ensure privacy in the software I create, I face some challenges. One challenge is that these tools may not fully understand privacy rules or the specific needs of the software. Privacy requirements can be complex, and the tools may struggle to implement customized privacy controls” – P39 (Developer, full-stack, 6 years exp.)
IV-C2 Poor output quality (9.80%; n=5)
Many developers experienced generation of buggy code, code which had discrepancies, and ended up with code that did not work properly. They also said that they found that the coding style is different from humans.
“I have encountered some challenges thus far - incorporating nuances of different privacy regulations across jurisdictions (different states and territories across Australia) can lead to discrepancies in the generated code” – P28 (Developer, back-end, desktop or enterprise applications, 20 years exp.)
IV-C3 Lack of trustworthiness - Distrust and use of alternatives to ensure privacy (64.71%; n=33)
Most of our surveyed developers said they do not trust AI-based programming assistants with sensitive data, and it is crucial to use them for this task very cautiously. They said that AI is not self-sufficient for privacy requirements, therefore, human intervention is necessary. Therefore, apart from writing manual code they also use other means such as generating code using example data and then amending it or relying on more reliable sources or colleagues.
“At the moment I still don’t blindly trust using these systems to write sensitive parts of the software, I prefer to rely on more reliable sources or the experience of colleagues who know better than me how to manage confidential information in the software we produce” – P19 (Developer, front-end, 4 years exp.)
IV-C4 Extra effort required (35.29%; n=18)
@@
However, many of these developers did believe that it was their responsibility to ensure that privacy requirements are met. Therefore, it is necessary to thoroughly review the code and perform necessities such as impact assessments.
“…developers still bear the responsibility for ensuring privacy requirements are met. It is essential to review and test the generated code thoroughly, conduct privacy impact assessments, and have appropriate mechanisms in place to address any privacy concerns or issues that may arise” – P1 (Developer, full-stack, 3 years exp.)
Summary: Current AI tools are only partially useful for generating code ensuring data privacy. The task characteristics and the technology characteristics misalign, hence in actual use many developers tend not to trust AI-based programming assistants for ensuring data privacy. The ones who do use it put extra effort into double-checking the code, which impacts their performance. Others use alternative methods to avoid using AI-based programming assistants.
IV-D Suggestions for Improvement of AI-based Programming Assistants for Privacy-related Code Generation (answer to RQ3)
@@
IV-D1 Understand Diverse Data Privacy Laws (24.75%; n=14)
@@
The most common suggestion from our developers is to train the AI to better ensure data protection in generated code. They mentioned that regular and more complete training and testing are necessary. They said that this training needs to be done using reliable and verified code.
“Perhaps these tools should be trained on reliable and verified code to make them generate code of a high enough quality to be used in corporate contexts when it comes to meeting privacy standards” – P17 (Developer, full-stack, 3 years exp.)
Further, they mentioned that AI needs to have bigger contextual thinking and robust algorithms that understand data privacy laws in different countries and regions which are significantly different [28].
IV-D2 Self Testing (7.84%; n=4)
@@
Even experienced developers make mistakes with privacy-related code [29]. Our developers suggested AI-based programming assistants should be more responsible and “self-test” the code provided. This would include checking that all privacy requirements are met by it before providing developers with the final code. This is important to generate better code which again will increase the safety of the code and usage of the tools.
“I think the generated code should undergo a test after generating to ensure the privacy requirements are met” – P39 (Developer, full-stack, 6 years exp.)
IV-D3 Explainable Output (3.92%; n=2)
@@
Current AI-based programming assistants fail to clearly explain rationale behind code produced. They should generate code with attention mechanisms and comments that explain to privacy-related code produced, including the list of the resources it has used.
“Techniques such as attention mechanisms or explainable AI methods can help developers identify potential privacy risks or vulnerabilities in the generated code” – P45 (Developer, back-end, front-end, mobile, 12 years exp.)
IV-D4 Improve Transparency (11.76%; n=6)
@@
Our developers said that the transparency of the owning company’s privacy policies and practices including data usage need to be improved. This is highly beneficial for AI code gen SaaS companies as building trust with the users will help in increasing the usage.
“We need better guarantee about the security of the tool itself, and the use of data by the company that owns it” – P25 (Developer, back-end, 10 years exp.)
IV-D5 Follow Standards (7.84%; n=4)
@@
From a governance point of view, some developers said that it is necessary to define a set of privacy requirements for the AI to follow. They said that following a set of lawful guidelines is necessary [30].
“The world needs to decide on what privacy requirements we are supposed to follow for current AI to work for it” – P12 (Developer, back-end, 3 years exp.)
IV-D6 Compliance (1.96%; n=1)
@@
One important suggestion was that AI-based programming assistants should show that they meet various compliance requirements [31]. The companies should audit their AI tools by an external party to certify them. Such certifications will bring better trustworthiness to the tools and hence to the code they generate and the software solutions they help to produce.
“Unfortunately there is no way from AI tool user to ensure at 100% that privacy requirements are met by AI tool, so users needs to trust the company that releases the AI tool. One way for having a certain level of ”privacy requirement ensuring” is to audit the AI tool by an external company, who certifies that privacy requirements are met” – P6 (Developer, full-stack, 10 years exp.)
IV-D7 Ease of Use (13.74%; n=7)

### v2

# How Are We Doing With Using AI-Based Programming Assistants For Privacy-Related Code Generation? The Developers’ Experience

@@
Abstract.
With generative AI becoming widespread, the existence of AI-based programming assistants for developers is no surprise. Developers increasingly use them for their work, including generating code to fulfil the data protection requirements (privacy) of the apps they build. We wanted to know if the reality is the same as expectations of AI-based programming assistants when trying to fulfil software privacy requirements, and the challenges developers face when using AI-based programming assistants and how these can be improved. To this end, we conducted a survey with 51 professional developers worldwide. We found that AI-based programming assistants need to be improved in order for developers to better trust them with generating code that ensures privacy. In this paper, we provide some recommendations including model and system-level improvements and some key further research directions to improve AI-based programming assistants for developing secure code.
@@ Abstract.
conference: The 29th International Conference on Evaluation and Assessment in Software Engineering; 17–20 June, 2025; Istanbul, Turkeycopyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/2018/06ccs: Software and its engineering Software notations and toolsccs: Software and its engineering Automatic programming
1. Introduction
@@
Given the capacity of generating human-like text using the large corpus of data the large language models (LLMs) were trained on, the popularity of using LLMs in applications of different domains is rising exponentially. Codex – OpenAI’s LLM model used to generate code in GitHub’s Copilot – claims that it is “proficient in more than a dozen programming languages, Codex can now interpret simple commands in natural language and execute them on the user’s behalf—making it possible to build a natural language interface to existing applications” [^1] https://openai.com/index/openai-codex/. Similar to Copilot, other LLM-driven code generators (AI-based programming assistants) such as Amazon Q Developer (Amazon Code Whisperer), Tabnine, Kite etc., are widely available for the developer community to use. These AI-based programming assistants claim that they enhance productivity, which is key in software development [^2] https://aws.amazon.com/q/developer/, https://github.com/features/copilot and some research has found the same results (Imai, 2022; Peng et al., 2023). However, AI-based programming assistants struggle with complex tasks (Tian et al., 2023), the accuracy of the code of these AI-based programming assistants depends on the programming language and the task (Nguyen and Nadi, 2022; Yetistiren et al., 2022; Moradi Dakhel et al., 2023), with longer details resulting in poor results (Cámara et al., 2023), and developers may spend more time reviewing the code that AI-based programming assistants produced (Bird et al., 2023). The developers also face challenges with understanding, editing, and debugging the code generated by the tools (Vaithilingam et al., 2022; Liang et al., 2023), resulting in inefficiency (Mozannar et al., 2022).
GDPR defines data protection (“privacy” as used in this paper), as “keeping data safe from unauthorised access” (General Data Protection Regulation, [n. d.]). Privacy engineering encourages embedding privacy into systems. For example, designing clear privacy controls (on the user-facing side), determining the best methods for anonymisation, and inspecting code before deployment to assess privacy risks. Proactive measures of privacy engineering include data minimisation and retention (Senarath and Arachchilage, 2019).
@@
For a business to leverage the data it has while ensuring the privacy of its users’ personal data, privacy-enhancing technologies (PETs) can be used. PETs are important due to the necessity of following data protection laws such as GDPR and CCPA, requirements of testing data by third parties, and protecting the business from tarnishing its reputation from privacy breaches. However, developers struggle with embedding privacy into software systems (Anthonysamy et al., 2017; Farhadi et al., 2019; Notario et al., 2015). They find it difficult to relate privacy requirements to privacy techniques, they see privacy in contradiction with system requirements, developers personal and peers’ opinions affect how they design, and they lack privacy knowledge (Senarath and Arachchilage, 2018a). When developers are asked to consider privacy in application designs, developers intuitively focus on technical aspects, but not user privacy expectations. Developers are also observed to have a reduced level of privacy expectations compared to users (Senarath and Arachchilage, 2018b). This leads to several critical issues. For example, permission-related issues frequently happen in Android apps (Scoccia et al., 2019). Having tools such as AI-based programming assistants comes in handy for developers in all the above-mentioned cases. Therefore, the developers may want to use AI-based programming assistants to generate code to meet their software’s privacy requirements.
@@
Nevertheless, the widespread uptake of AI-based programming assistants to implement systems – including data privacy-related code – raises the potential of their use for more secure code construction but, on the other hand, also introduces unexpected and unintended code vulnerabilities (Gupta et al., 2023). To better understand the practical experience of using AI-based programming assistants for privacy-related code generation, we carried out a survey [^3] Approved by Monash Human Research Ethics Committee. Project ID: 37904. of 51 software developers worldwide. We asked the developers about their experience with AI-based programming assistants for privacy-related code generation. Our research yielded knowledge and insights on AI-based programming assistants, including the difference between the expectation and reality of AI-based programming vis-à-vis meeting privacy-related requirements in AI-generated code and the critical need for improvements in the AI-based programming assistants for developers to trust them with generating code for ensuring privacy.
2. Research Questions
We wanted to answer the following research questions in this work:
@@ 2. Research Questions
    Is the reality the same as the expectation of AI-based programming assistants when it comes to ensuring privacy requirements are met in the code? Privacy is a crucial non-functional requirement in software development. We deep-dived into this aspect on the subject of the code generated by AI-based programming assistants. As the first step, we evaluated if the reality is similar or different to developers’ expectations (Section 4.2).
@@
    What are the challenges with generating AI-based programming assistants to ensure meeting privacy requirements? We wanted to explore the challenges concerning privacy-ensuring code as identifying the challenges is essential for improving the AI-based programming assistants (Section 4.3).
@@
    How can AI-based programming assistants be improved? Developers are the users of AI-based programming assistants. It is critical to get their feedback on improvements to these programming assistants based on their experiences. We used this question to investigate potential improvements in AI-based programming assistants (Section 4.4).
3. Our Approach – The Developer Survey
@@
Survey Questionnaire Development. To answer the research questions, we conducted a developer survey. We designed our survey with both open-ended and closed-ended questions with a mapping against our research questions. For example, to answer RQ1, we used both open-ended and closed-ended questions. For the complete survey questionnaire (this includes questions beyond the scope of this paper), please check our online replication package  [^4] https://github.com/kashumi-m/AICodeGensPrivacy. The questions used to collect data on company size, developers’ roles, and employment status reported in this paper were adapted from the Stack Overflow developer survey [^5] https://survey.stackoverflow.co/2022. After we developed the questionnaire, we piloted it with the help of two postdoctoral researchers who worked in the software industry in the past at the HumaniSE Lab at Monash University. They gave feedback on completion time, and the flow of the questions. Upon their feedback, we moved the demographic questions to the end of the questionnaire, which were at the beginning earlier.
@@
Data Collection and Analysis. After finalising the questionnaire, we hosted the survey on Qualtrics and recruited developers who work in the IT industry via the recruitment platform Prolific. We spent 6 Sterling Pounds on each participant. The participant spent around 15-20 minutes to fill out the questionnaire. We recruited 51 participants and we were able to generate rich insights from the data we collected. The details of the participants are given in the next section. To analyse the data we collected through open-ended (qualitative) and closed-ended (quantitative) questions, we used open coding and constant comparison (Strauss and Corbin, 1998) – interpreting the data using small chunks of words (codes) and categorising them by comparison, and with descriptive statistics.
4. Findings
4.1. Survey Participants
4.1.1. Demographics of the Participants
@@
The majority of the developers who participated in our study were men (76.47%; n=39) and the participants resided around the world. Most participants were from companies with 101–500 employees (17.65%; n=9). The majority were from the IT industry (78.43%; n=40). Most of our participants were full-time employees (88.84%; n=45) and the majority worked hybrid – some remote and some in-person (52.94%; n=27). The participants had a median of 9 years of experience in the software industry, 7 years of professional coding experience, and 6 years of agile experience.
4.1.2. Who Used AI-based Programming Assistants the Most and Who Used Them the Least
@@
While our participants played multiple roles in their jobs, the majority were full-stack developers (37.25%; n=19). The majority of the developers used AI-based programming assistants 1–2 times a day (54.90%; n=28). From that, most were full-stack developers (21.57%; n=11). Out of the participants who never used AI-based programming assistants (19.61%; n=10) were the back-end developers (7.84%; n=4). ChatGPT was the most commonly used (61.54%; n=32) AI-based programming assistant by our developers.
@@
4.2. RQ1 – Expectation vs Reality of AI-based Programming Assistants for Privacy-related Code Generation
@@
When first using an AI-based programming assistant (tool), the majority of the developers (64.71%; n=33) expected that the tool will generate code ensuring that key data privacy requirements are met (Fig. 1). However, as they moved forward with using the tool, reality became somewhat different from what they had expected. The majority (39.22%; n=20) found that AI-based programming assistants sometimes generate code ensuring privacy requirements without requesting it to meet the requirement (Fig. 2). However, they had to always (52.94%; n=27) put extra effort into double-checking the generated code to see if privacy requirements were met (Fig. 3). Therefore, unsurprisingly, a large number of the developers (43.14%; n=22) never used AI-based programming assistants to generate code for privacy requirements (Fig. 5), and many of them (41.18%, n=21) wrote code manually most of the time to meet privacy requirements (Fig. 4), even though they used AI-based programming assistants during software development.
@@
Figure 1. Developers expected that the AI-based programming assistant will ensure meeting the privacy requirements.
@@
Figure 2. Developers found that AI-based programming assistants sometimes generate code meeting privacy requirements.
@@
Figure 3. Developers always put extra effort to double-check the code generated by the AI tool to ensure the privacy requirements were met.
@@
Figure 4. Most of the time, the majority of the developers wrote code manually to ensure privacy requirements were met.
@@
Figure 5. Majority of the developers never used AI-based programming assistants to generate code to meet privacy requirements.
@@
4.3. RQ2 – Key Challenges in Generating Code Ensuring Privacy using AI-based Programming Assistants
@@
4.3.1. Unfulfilled task - not generating privacy-preserving code (19.61%; n=10)
Our developers mentioned that current AI-based programming assistants lack deep comprehension of and code generation meeting diverse privacy regulations. They said that AI-based programming assistants may not be able to keep up with changing privacy policies, they may not fully understand privacy rules or specific needs. Hence, the AI-based programming assistants may struggle to implement customised privacy controls: “When I use AI-code generators to ensure privacy in the software I create, I face some challenges. One challenge is that these tools may not fully understand privacy rules or the specific needs of the software. Privacy requirements can be complex, and the tools may struggle to implement customized privacy controls” – P39 (Developer, full-stack, 6 years exp.)
4.3.2. Poor output quality (9.80%; n=5)
Many developers experienced generation of buggy code, code which had discrepancies, and ended up with code that did not work properly. They also said that they found that the coding style is different from humans: “I have encountered some challenges thus far - incorporating nuances of different privacy regulations across jurisdictions (different states and territories across Australia) can lead to discrepancies in the generated code” – P28 (Developer, back-end, desktop or enterprise applications, 20 years exp.)
4.3.3. Lack of trustworthiness - Distrust and use of alternatives to ensure privacy (64.71%; n=33)
Most of our surveyed developers said they do not trust AI-based programming assistants with sensitive data, and it is crucial to use them for this task very cautiously. They said that AI is not self-sufficient for privacy requirements, therefore, human intervention is necessary. Therefore, apart from writing manual code they also use other means such as generating code using example data and then amending it or relying on more reliable sources or colleagues: “At the moment I still don’t blindly trust using these systems to write sensitive parts of the software, I prefer to rely on more reliable sources or the experience of colleagues who know better than me how to manage confidential information in the software we produce” – P19 (Developer, front-end, 4 years exp.)
4.3.4. Extra effort required (35.29%; n=18)
@@
However, many of these developers did believe that it was their responsibility to ensure that privacy requirements are met. Therefore, it is necessary to thoroughly review the code and perform necessities such as impact assessments: “…developers still bear the responsibility for ensuring privacy requirements are met. It is essential to review and test the generated code thoroughly, conduct privacy impact assessments, and have appropriate mechanisms in place to address any privacy concerns or issues that may arise” – P1 (Developer, full-stack, 3 years exp.)
4.4. RQ3 – Suggestions for Improvements to AI-based Programming Assistants for Privacy-related Code Generation
@@
4.4.1. Understand Diverse Data Privacy Laws (24.75%; n=14)
@@
The most common suggestion from our developers is to train the AI to better ensure data protection in generated code. They mentioned that regular and more complete training and testing are necessary. They said that this training needs to be done using reliable and verified code: “Perhaps these tools should be trained on reliable and verified code to make them generate code of a high enough quality to be used in corporate contexts when it comes to meeting privacy standards” – P17 (Developer, full-stack, 3 years exp.). Further, they mentioned that AI needs to have bigger contextual thinking and robust algorithms that understand data privacy laws in different countries and regions which are significantly different (Baumer et al., 2004).
4.4.2. Self Testing (7.84%; n=4)
@@
Even experienced developers make mistakes with privacy-related code (Oliveira et al., 2018). Our developers suggested AI-based programming assistants should be more responsible and “self-test” the code provided. This would include checking that all privacy requirements are met by it before providing developers with the final code. This is important to generate better code which again will increase the safety of the code and usage of the tools: “I think the generated code should undergo a test after generating to ensure the privacy requirements are met” – P39 (Developer, full-stack, 6 years exp.)
4.4.3. Explainable Output (3.92%; n=2)
@@
Current AI-based programming assistants fail to clearly explain rationale behind code produced. They should generate code with attention mechanisms and comments that explain to privacy-related code produced, including the list of the resources it has used: “Techniques such as attention mechanisms or explainable AI methods can help developers identify potential privacy risks or vulnerabilities in the generated code” – P45 (Developer, back-end, front-end, mobile, 12 years exp.)
4.4.4. Improve Transparency (11.76%; n=6)
@@
Our developers said that the transparency of the owning company’s privacy policies and practices including data usage need to be improved. This is highly beneficial for AI code gen SaaS companies as building trust with the users will help in increasing the usage: “We need better guarantee about the security of the tool itself, and the use of data by the company that owns it” – P25 (Developer, back-end, 10 years exp.)
4.4.5. Follow Standards (7.84%; n=4)
@@
From a governance point of view, some developers said that it is necessary to define a set of privacy requirements for the AI to follow. They said that following a set of lawful guidelines is necessary (Gasiba et al., 2020): “The world needs to decide on what privacy requirements we are supposed to follow for current AI to work for it” – P12 (Developer, back-end, 3 years exp.)
4.4.6. Compliance (1.96%; n=1)
@@
One important suggestion was that AI-based programming assistants should show that they meet various compliance requirements (Klymenko et al., 2022). The companies should audit their AI tools by an external party to certify them. Such certifications will bring better trustworthiness to the tools and hence to the code they generate and the software solutions they help to produce: “Unfortunately there is no way from AI tool user to ensure at 100% that privacy requirements are met by AI tool, so users needs to trust the company that releases the AI tool. One way for having a certain level of ”privacy requirement ensuring” is to audit the AI tool by an external company, who certifies that privacy requirements are met” – P6 (Developer, full-stack, 10 years exp.)
4.4.7. Ease of Use (13.74%; n=7)